What is the molar solubility of CaF2 in a solution containing 0.100 M NaF? (Ksp for...

  1. What is the molar solubility of CaF2 in a solution containing 0.100 M NaF? (Ksp for CaF2 is 1.46 x 10-10?
Answer #1

First we determine the dissociation that occurs:

CaF2 = Ca + 2 + 2F-

In the solution there is already a concentration of 0.1 M of ions of F-, then, in the equilibrium:

[CaF2] = X

[Ca + 2] = X

[F-] = 0.1 + 2X

Substituting the Kps expression:

Kps = [Ca + 2] * [F -] ^ 2 = X * (0.1 + 2X) ^ 2

simplifying:

4X ^ 3 + 0.4X ^ 2 + 0.01X - 1.46x10 ^ -10 = 0

clearing:

X = Molar Solubility = 1.46x10 ^ -8 M